what is the area of a circle whose diameter is 3 centimeters?
Answer by Waltham(3)

Diameter=3cm
Radius=Diameter/2=1.5cm
Area=πr^2
substitute pi for 22/7
and radius, r for 1.5 and find the area
answer=1.5^2 times 22/7=49.5/7 or about 7